# Inferring Scientific Cross-Document Coreference and Hierarchy with Definition-Augmented Relational Reasoning

### v2

@@ 4.2 Results
In addition to reporting overall performance (see Table 8), we focus particular attention on the "hard" data subsets provided by Cattan et al. (2021b), comprised of the 10% (hard 10) or 20% (hard 20) topics (collections of candidate mentions) with lowest scores as ranked according to CoNLL F1 score by using Levenshtein edit distance as a baseline in agglomerative clustering. The subsets are particularly challenging, exhibiting high lexical diversity and ambiguity, and are thus of most interest; e.g., the best model reported by Cattan et al. (2021b), achieved only 64.34 F1 for coreference CoNLL F1 on the hard 10 subset, while achieving 76.46 F1 on the full test set.
We report our results in Table 1 (hard 10) along with Tables 10, 11, 12, 13, 14, 15 covering the hard 20 and general test subsets in the Appendix. ICL few-shot results are averaged across 5 runs as described earlier. We first observe that, not very surprisingly, fine-tuning of recent LLMs leads to large gains over the smaller Longformer model used originally in Cattan et al. (2021b). Fine-tuning also blows out of the water concurrent results for SciCo CDCR in the few-shot setting with larger variants from the Mistral family (see Table 3 in Lior et al. (2024)), highlighting the value of fine-tuning in this task. More interestingly, our contextualized definitions led to strong improvements, for both LLM fine tuning and ICL. Our SciCo-RadarRelational model, which consists of fine-tuning Mistral-7B on SciCo augmented by relational definitions, achieved the highest results in every evaluation category, as exemplified in Table 1. Generally, relational definitions outperformed singleton definitions, by carrying more nuanced information about relationships that the model might miss with the more generic singleton definitions (see Section 5 for more analysis). ICL with larger LLMs OpenAI (2024), including prompt optimization with DSPy, unperformed smaller fine-tuned models in our experiments, though future work may explore this further with more models and in-context examples.
As seen in Tables 13, 10, augmenting Longformer with our contextual definitions also resulted in significant enhancements in both hierarchy and coreference scores, by up to 3.38 and 1.72 respectively. Even more significant improvements were observed within the hard 10 and hard 20 subsets, particularly in hierarchy and hierarchy 50% scores, achieving improvements of 2.84 and 4.16 in the hard 20 subset (Tables 14, 11). With Mistral-7B, incorporating definitions consistently improved all metrics. As seen In Table 1, incorporating definitions notably boosted
Mistral’s hierarchy F1 and F1-50% scores by up to 4.5 and 6.1 using singleton definitions, and by 6.78 and 8.63 with relational definitions.
We also report additional coreference and hierarchy precision-recall metrics in Tables 13, 14 and 10, 11 in the Appendix. For coreference, the F1 gains with relational definitions come from a significant boost in recall with only a relatively small dip in precision. For hierarchy, we see a similar trend where both singleton and relational definitions improve recall while maintaining stable precision.
Finally, to assess the stability of ICL results with random few-shot examples or examples optimized by random search in DSPy, we conduct five runs and report average results with standard errors (Table 3).999The SciCo dataset contains 69,682 mention pairs in the hard 10 subset, along with extensive context and definitions per mention; running our ICL experiments on this subset cost $500. The full test set contains 564,352 mention pairs. We find that the effects are overall stable, with definitions helping boost results across both the few-shot and DSPy experiments;101010We also experimented with several runs of fine-tuning, finding negligible variance across seeds. interestingly, we do not observe gains from using DSPy in our experiments.
| Model | Coreference | Hierarchy | Path | |
| CoNLL F1 | F1 | F1-50% | Ratio | |
| GPT 4o-mini | 51.28 ± 2.21 | 40.22 ± 1.70 | 16.95 ± 1.62 | 31.04 ± 2.23 |
| GPT 4o-miniDSPy | 50.71 ± 0.66 | 42.23 ± 0.94 | 14.74 ± 1.36 | 27.41 ± 1.29 |
@@
As an additional baseline and ablation, we use the powerful and more costly GPT-4o to generate definitions by providing the model with the mention and its surrounding context but no additional retrieval, using only the model’s vast knowledge embedded in its parameter space Ye et al. (2023); Bubeck et al. (2023). We find that after training Mistral-7B on these definitions, a big gap in performance emerged in favor of retrieval-based definitions in terms of hierarchy predictions (§ Table 7), along with a smaller but noticeable advantage in coreference scores, too (§ Table 15). Upon reviewing the generated definitions, we observed that GPT-4o often generated definitions that were more generic and tended to be more fixated on irrelevant context details, while retrieval-based definitions naturally assisted the definition generation to be both more detailed and specific while zooming out and abstracting-away irrelevant context (Table 5; §5 for more analysis).
@@ 5.1 When Definitions Generally Help
We begin with providing qualitative evidence to support the fundamental hypothesis in this paper, showing that incorporating dynamic definitions significantly improves model performance over the non-definition approach. We discuss two main ways in which we qualitatively observe definitions broadly help, before diving deeper into specific types of definitions and their various trade-offs.
Lexical ambiguity, diversity, and fine-grained hierarchy
Lexical ambiguity and diversity, and fine-grained concept hierarchies are at the heart of the challenge for the H-CDCR task in complex technical domains. Our findings confirm our hypothesis that definitions provide useful context that helps models understand lexical and semantic nuances for resolving ambiguity and diversity, detecting shared meanings and aspects that are not obvious from the text alone. For instance (see also Table 2, row 2), the definition of online marketing operation describes it as encompassing various digital strategies, including targeted ads, clarifying that the former subsumes the latter. Similarly, defining targeted ads as a specific type of online advertisement tailored to specific demographics makes the hierarchical relationship clear. As another example, demonstrating nuanced lexical ambiguity (see also Table 2, row 1), the definitions for network cyber-security and cyber-safety awareness make it clear that one mention focuses on technical measures to secure network infrastructure, while the other emphasizes educating individuals about online safety practices. These definitions help the model understand that, despite surface similarities, the mentions do not belong to the same conceptual cluster or share a hierarchical link.
@@
We observe that, unsurprisingly, vague or ambiguous contexts can make it hard to clarify whether two different mentions are co-referential, hierarchical, or unrelated. Additionally, a mention might be a minor part of a broader discussion where the surrounding text focuses on other concepts, potentially injecting irrelevant context into the model’s decision on the cross-document relation. We observe that definitions generally help mitigate these challenges by providing detailed explanations that highlight similarities or shared concepts that extend beyond the surface-level context. For example, conversational speech synthesis and conversational speech processing tasks (see Table 2, row 3) where the first mention is part of a broader discussion on F0 control and speaking attitudes, and the second mention involves a focus on privacy-preserving features in speech processing.
@@ 5.2 Impact of Relational Definitions
Relational definitions can sometimes emphasize similarities or overlaps that are not relevant to the actual relationship between the mentions. Relational definitions are created for pairs that have been classified as connected, and while they are adept at highlighting differences, as seen in previous examples, they may still overemphasize similarities. For example, the relational definitions for Transformer model and Equivariant Transformers (ETs) (see Table 6 in the Appendix) incorrectly identifies a hierarchical relationship, while the mentions refer to different architecture families.
| Context and Prediction | Singleton Def | GPT4o Def |
Ground Truth: |
Correct Prediction: |
Wrong Prediction: |
Ground Truth: |
Correct Prediction: |
Wrong Prediction: |
@@
Finally, we qualitatively examine how retrieval definitions differ to GPT4o-based definitions. We find that generally, retrieval generates more detailed and comprehensive definitions covering more diverse concept aspects, that allow detecting when one concept is more general than the other; conversely, GPT4o tends to generate more vanilla, less rich definitions (e.g., see Table 5, row 1). Additionally GPT4o definitions tend more often to fixate and over-rely on context, making the definitions focused on too narrow aspects of the context which do not allow to zoom out and detect abstractions and subsumptions (see example in Table 5, row 2). These findings help explain why retrieval definitions appear to add several points in Hierarchy F1 vs. GPT4o definitions (Table 15 in the Appendix).
@@ 6 Related Work
Our task is closely related to the cross-document co-reference resolution task [Yang et al. (2015), Cattan et al. (2021a)]. In scientific domains, CDCR presents unique challenges due to the abstract nature of technical concepts and the need to differentiate between closely related terms at varying levels of specificity Cattan et al. (2021b). Recent advancements in coreference resolution have been reached with LLMs. As demonstrated by [Le and Ritter (2023), Gan et al. (2024)], LLMs like GPT-4, InstructGPT and Llama variants, when employed with prompt-based techniques, have shown strong performance in co-reference resolution tasks in the more popular within-document setting. In our approach, we explore LLMs in the challenging cross-document setting, in the scientific domain, and include also cross-document hierarchy detection. We use both in-context learning and fine-tuning with dynamically generated (relational) definitions to improve CDCR performance.
@@
Munnangi et al. (2024) recently enhanced biomedical NER capabilities of LLMs in the few-shot setting by incorporating static definitions of canonicalized entities primarily from the UMLS knowledge base Bodenreider (2004). In our method, we generate dynamic singleton and relational definitions specifically tailored to the unique context of the mentions in scientific texts, by retrieving relevant literature. Our definitions are designed to enhance understanding of cross-document relationships. Xin et al. (2024) recently proposed a method that uses LLMs as context augmenters to enhance entity linking (EL) with mention descriptions which are integrated into traditional EL models to improve linkage with canonical entities in general-domain KBs.
@@
There has been scarce work we are aware of on generating relational concept descriptions. Murthy et al. (2022) introduced a system named ACCoRD for generating diverse descriptions of how scientific concepts are related in order to provide varied perspectives on scientific concepts. The focus of ACCoRD was to build a small-scale user-facing system demonstration; unlike ACCoRD, our work uses relational definitions to directly improve NLP tasks such as CDCR and hierarchy inference.
@@ 7 Conclusion
We presented SciCo-Radar, a novel approach based on dynamic definition generation and relational reasoning for enhancing the ability of Large Language Models (LLMs) to infer cross-document coreference and hierarchy relationships in scientific papers. Our method employs context-dependent and relational definitions generated through literature retrieval, along with a re-ranking based scheme to avoid combinatorial explosion across all potential mention pairs. SciCo-Radar has shown substantial improvements in both fine-tuning and in-context learning settings. We conducted in-depth analyses into the quality of generated definitions and failure modes which could be valuable to explore in future work. Future research may focus on improving the generation and utilization of dynamic concept definitions, as well as improving the relational reasoning capabilities of LLMs when handling nuanced, fine-grained scientific concepts. Additionally, exploring ways to further optimize the efficiency of our two-stage re-ranking approach could lead to even more scalable solutions for large-scale scientific applications. Finally, it is exciting to consider broader applications of our method, across more cross-document reasoning tasks.
8 Limitations
We note a few limitations, which could be explored in future work. A basic limitation is that the quality of automated contextual definitions relies on the performance of the underlying language models, which may produce inaccurate or overly general descriptions in some cases. Using large-scale state-of-art LLMs for generating defintions can incur prohibitive costs when working at the scale of scientific literature; in this work, we use smaller but still high-performance models that in practice produce mostly high-quality definitions. In addition, generating relational definitions for mention pairs can be computationally intensive and potentially lead to combinatorial explosion with infeasible costs. In our work, we thus do not apply relational definitions across all possible mention pairs, instead developing a re-ranking method with relational definitions. An additional limitation relates to the SciCo dataset, which despite its relatively large scale (over 3x larger than ECB+ Cybulska and Vossen (2014)), still does not come close to capturing the full breadth of scientific literature and its complexity (e.g., it only focuses on computer science papers, and the number of challenging/“hard” topics could in principle be made dramatically larger in future benchmarks for the fundamental task of H-CDCR).
@@ Appendix A Appendix
A.1 ICL
@@
We select 200 pairs for training and 60 pairs for development. To ensure balanced evaluation, we drew a random subset of examples, ensuring an equal number of labeled examples across the different relations. Within our DSPy program we incorporate chain-of-thought (CoT) reasoning and BootstrapFewShotWithRandomSearch optimizer 111111https://dspy-docs.vercel.app/docs/building-blocks/optimizers. In CoT, a model is instructed to reason step by step to get to its answer, which can improve performance Wei et al. (2022). BootstrapFewShotWithRandomSearch is an enhanced version of the BootstrapFewShot teleprompter 121212https://dspy-docs.vercel.app/docs/deep-dive/teleprompter/bootstrap-fewshot, which uses a teacher (GPT4 OpenAI (2023)) - student (GPT4o-mini) approach to generate few-shot demonstrations by sampling from the training data. During the training phase, the teacher model simulates potential outputs and rationales for labeled instances. The teacher model iteratively generates few-shot examples by attempting to predict the correct outputs based on the training data, learning from its own successes and refining also the generated rationale over time. This process is repeated multiple times, with a random search conducted over the generated demonstrations. The chosen program is the one that achieves the highest performance on the dev set. During each iteration on the test set, the program receives few-shot examples along with their bootstrapped rationales, assisting it in forming better rationales, thereby resulting in improved predictions.

### v3

@@ 4.2 Results
As an additional baseline and ablation, we use the powerful and more costly GPT-4o to generate definitions by providing the model with the mention and its surrounding context but no additional retrieval, using only the model’s vast knowledge embedded in its parameter space Ye et al. (2023); Bubeck et al. (2023). We find that after training Mistral-7B on these definitions, a big gap in performance emerged in favor of retrieval-based definitions in terms of hierarchy predictions (Table 12), along with a smaller but noticeable advantage in coreference scores, too (Table 19). Upon reviewing the generated definitions, we observed that GPT-4o often generated definitions that were more generic and tended to be more fixated on irrelevant context details, while retrieval-based definitions naturally assisted the definition generation to be both more detailed and specific while zooming out and abstracting-away irrelevant context. For example, in Table 11, the GPT-4o definition of “representation learning” is too focused on a specific application involving embodied agents and cognitive phenomena, while the singleton definition abstracts away from these details and instead highlights the broader machine learning context, the kind of generalization needed to resolve this hierarchical relationship (Table 11 in the Appendix; §5 for more analysis).
@@ 5.1 When Definitions Generally Help
We begin with providing qualitative evidence to support the fundamental hypothesis in this paper, showing that incorporating dynamic definitions substantially improves model performance over the non-definition approach. We discuss two main ways in which we qualitatively observe definitions broadly help, before diving deeper into specific types of definitions and their various trade-offs.
Lexical ambiguity, different surface forms, and fine-grained hierarchy
Lexical ambiguity, different surface forms, and fine-grained concept hierarchies are at the heart of the challenge for the H-CDCR task in complex technical domains. Our findings confirm our hypothesis that definitions provide useful context that helps models understand lexical and semantic nuances for resolving the true relationships of mentions, or lack there of, by detecting shared meanings and aspects that are not obvious from the text alone. For instance (see also Table 5, row 2), the definition of online marketing operation describes it as encompassing various digital strategies, including targeted ads, clarifying that the former subsumes the latter. Similarly, defining targeted ads as a specific type of online advertisement tailored to specific demographics makes the hierarchical relationship clear. As another example, demonstrating lexical ambiguity (see also Table 5, row 1), the definitions for network cyber-security and cyber-safety awareness make it clear that one mention focuses on technical measures to secure network infrastructure, while the other emphasizes educating individuals about online safety practices. These definitions help the model understand that, despite surface similarities, the mentions do not belong to the same conceptual cluster or share a hierarchical link.
@@
We observe that, unsurprisingly, vague or ambiguous contexts can make it hard to clarify whether two different mentions are co-referential, hierarchical, or unrelated. Additionally, a mention might be a minor part of a broader discussion where the surrounding text focuses on other concepts, potentially injecting irrelevant context into the model’s decision on the cross-document relation. We observe that definitions generally help mitigate these challenges by providing detailed explanations that highlight similarities or shared concepts that extend beyond the surface-level context. For example, conversational speech synthesis and conversational speech processing tasks (see Table 5, row 3) where the first mention is part of a broader discussion on F0 control and speaking attitudes, and the second mention involves a focus on privacy-preserving features in speech processing. We note that this type of context can affect the quality of definitions as well; however, the retrieved global literature information may help mitigate problems in the local context, such that the definition generation model can often successfully generate a useful enough definition to assist the classification model.
@@ 5.2 Impact of Relational Definitions
Relational definitions may sometimes emphasize similarities or overlaps irrelevant to the actual relation between mentions. For example, the relational definitions for Transformer model and Equivariant Transformers (ETs) (see Table 10 in the Appendix) incorrectly identifies a hierarchical relationship, while the mentions refer to different architecture families.
@@
Finally, we qualitatively examine how retrieval definitions differ to GPT4o-based definitions. We find that generally, retrieval generates more detailed and comprehensive definitions covering more diverse concept aspects, that allow detecting when one concept is more general than the other; conversely, GPT4o tends to generate more vanilla, less rich definitions (e.g., see Table 11, row 1 in the Appendix). Additionally GPT4o definitions tend more often to fixate and over-rely on context, making the definitions focused on too narrow aspects of the context which do not allow to zoom out and detect abstractions and subsumptions (see example in Table 11, row 2 in the Appendix). These findings help explain why retrieval definitions appear to add several points in Hierarchy F1 vs. GPT4o definitions (Table 19 in the Appendix).
@@ 6 Related Work
Our task is closely related to the cross-document co-reference resolution task Yang et al. (2015); Cattan et al. (2021a). In scientific domains, CDCR presents unique challenges due to the abstract nature of technical concepts and the need to differentiate between closely related terms at varying levels of specificity Cattan et al. (2021b). Ravenscroft et al. (2021) further extend this line of work by introducing a cross-document, cross-domain coreference resolution task, which addresses coreference resolution between scientific articles and related news reports. Their work emphasizes the linguistic complexity introduced by domain-specific variations, highlighting significant limitations of the existing CDCR methods. Finin et al. (2009) explored leveraging external structured knowledge through Wikitology, a knowledge base built from Wikipedia, DBpedia, and Freebase, demonstrating that incorporating structured semantic information substantially improves cross-document entity coreference performance. Recent advancements in coreference resolution have been reached with LLMs. As demonstrated by Le and Ritter (2023), Gan et al. (2024), LLMs like GPT-4, InstructGPT and Llama variants, when employed with prompt-based techniques, have shown strong performance in co-reference resolution tasks in the more popular within-document setting. Yang et al. (2022) analyzed GPT models ability to resolve coreference using QA-based prompting. They found that while those models can generate reasonable answers, their coreference resolution performance is inconsistent and highly sensitive to prompt phrasing, which highlights some limitations of LLMs in those domains. In our approach, we explore LLMs in the challenging cross-document setting, in the scientific domain, and include also cross-document hierarchy detection. We use both in-context learning and fine-tuning with dynamically generated (relational) definitions to improve CDCR performance.
@@
Earlier works leveraged external knowledge to improve text understanding. Al-Harbi et al. (2017) addressed lexical ambiguity in QA by incorporating WordNet-based definitions and domain-specific ontologies to refine word meanings. By integrating context knowledge, their approach improved word sense disambiguation, ensuring more precise retrieval in question answering. Yin and Roth (2018) similarly used dictionary definitions to improve hypernymy detection, addressing the limitations of distributional embeddings, which infer meaning based on co-occurrence patterns in large corpora. Although the usage of static definitions can be effective, they are often insufficient in our cross-document coreference and hierarchy task. Terms may have context dependent meanings that can vary across different papers, research fields, or even within the same document. In our approach, instead of relying on predefined dictionary definitions, we incorporate contextual definitions, generated with retrieved external knowledge in order to capture the nuanced semantic distinctions or similarities.
More recent works includes Munnangi et al. (2024) who enhanced biomedical NER capabilities of LLMs in the few-shot setting by incorporating static definitions of canonicalized entities primarily from the UMLS knowledge base Bodenreider (2004). In our method, we generate dynamic singleton and relational definitions specifically tailored to the unique context of the mentions in scientific texts, by retrieving relevant literature. Our definitions are designed to enhance understanding of cross-document relationships. Xin et al. (2024) recently proposed a method that uses LLMs as context augmenters to enhance entity linking (EL) with mention descriptions which are integrated into traditional EL models to improve linkage with canonical entities in general-domain KBs. Silva et al. (2020) introduce XTE (Explainable Text Entailment), an approach that explicitly leverages structured lexical definitions to improve semantic reasoning in RTE (Recognizing Textual Entailment). While their approach addresses entailment rather than cross-document coreference, it similarly demonstrates the benefit of augmenting neural models with structured concept definitions to enhance performance in relational reasoning tasks.
@@
There has been scarce work we are aware of on generating relational concept descriptions. Murthy et al. (2022) introduced a system named ACCoRD for generating diverse descriptions of how scientific concepts are related in order to provide varied perspectives on scientific concepts. The focus of ACCoRD was to build a small-scale user-facing system demonstration; unlike ACCoRD, our work uses relational definitions to directly improve NLP tasks such as CDCR and hierarchy inference. Brahman et al. (2021) generate natural language definitions in the form of rationales to explain defeasible inference, a type of nonmonotonic reasoning where new information can strengthen or weaken prior conclusions. These definitions include both local rationales for individual concepts and relational explanations of how updates alter entailment. They are generated by prompting a pretrained language model with a premise, hypothesis, and update, and trained via distant supervision without retrieval. While focused on entailment rather than coreference, their work similarly highlights the value of textual definitions for modeling conceptual relationships.
@@ 7 Conclusion
We presented SciCo-Radar, a novel approach based on dynamic definition generation and relational reasoning for enhancing the ability of Large Language Models (LLMs) to infer cross-document coreference and hierarchy relationships in scientific papers. Our method employs context-dependent and relational definitions generated through literature retrieval, along with a re-ranking based scheme to avoid combinatorial explosion across all potential mention pairs. SciCo-Radar has shown consistent improvements across both fine-tuning and in-context learning settings, with particularly strong gains on challenging subsets, and moderate improvements on the overall dataset. We showed that with our finetuning and dynamic definition integration approach, smaller open models can outperform much larger models. We also conducted in-depth analyses into the quality of generated definitions and failure modes which could be valuable to explore in future work. In addition, paired bootstrap tests provide further evidence that the improvements, particularly on the hardest subsets, are consistent across resampling.
Future research may focus on improving the generation and utilization of dynamic concept definitions, as well as improving the relational reasoning of LLMs when handling nuanced, fine-grained scientific concepts. Generating relational definitions can be computationally intensive at scale; we thus do not apply relational definitions across all possible mention pairs. Exploring ways to further optimize the efficiency of our re-ranking approach, or other approaches for explicit relational reasoning that avoid combinatorial explosion with infeasible costs, may lead to more scalable solutions for large-scale scientific applications. Another interesting direction is to investigate how the addition of structured metadata such as publication venue, field of study or year can help. For example, the term transformer may be interpreted as a neural model in a machine learning conference but as power equipment in an electrical engineering journal. Metadata signals could enhance both the generation of contextual definitions and the prediction of cross-document relations, particularly in ambiguous or long-tail cases. In addition, while orthogonal to our contribution, we note that the SciCo dataset (which is over 3 times larger than the common CDCR benchmark ECB+ Cybulska and Vossen (2014)) could be expanded by collecting more annotations of coreference clusters and hierarchies across many papers from different domains, to better capture the breadth and scale of literature. While our approach improves performance, particularly on harder examples, it comes with additional computational cost due to relational definition generation and fine-tuning. This cost also carries environmental impact especially in large-scale settings. Developing more efficient alternatives for relational reasoning and knowledge integration is not only important for scalability but also for reducing environmental footprint. We leave exploration of lower-cost yet effective variants as a promising direction for future work. Finally, it would be exciting to consider further applications of our method, in more general cross-document tasks.
Acknowledgments
We thank the reviewers and the action editor for their valuable feedback and suggestions. This research was supported by the Azrieli Foundation.
@@ Appendix A Appendix
A.1 Definition Generation technical details
For each definition generation iteration we retrieve the relevant papers, using mxbai-embed-large-v1 Lee et al. (2024); Li and Li (2024) as an embedding model with state-of-the-art performance. We use the Chroma DB101010https://github.com/chroma-core/chroma open-source vector database for fast retrieval using approximate nearest neighbor search. We rerank retrieved results with a cross-encoder (mxbai-rerank-large-v1111111mixedbread.ai/docs/reranking/models. Results are then finally fed into the definition generation model, for which we use Mixtral 8x7B, a mixture of experts LLM Jiang et al. (2024). We employ this method to produce singleton definitions for each individual mention and relational definitions (see Figures 4-5 for specific prompts and A.2 for hardware details and limitations).
A.2 Additional Hardware and LoRA Hyper-Parameter Details
Hardware Setup and Runtime
We employed 8 NVIDIA RTX A6000 48GB GPUs to generate both singleton and relational definitions through our RAG pipeline. Generating relational definitions only for the top-ranked mention pairs (§3) took roughly two weeks, while to generate relational definitions for the entire test set, we would need roughly 40 days. For Mistral-7B, Fine-tuning was conducted on 3 NVIDIA RTX A6000 48GB GPUs over 2 epochs. The training was carried out with a batch size of 2, 4 gradient accumulation steps and a maximum sequence length of 1664 tokens (these choices were made due to our academic resources). We used a default learning rate of 2e-5 with a weight decay of 0.001.
LoRA Configuration and Hyper-Parameters
We use Low-Rank Adaptation of LLM (LoRA), a parameter efficient fine-tuning method Hu et al. (2022) with the following parameters selected according to our resource constraints: lora_alpha: 16, lora_dropout: 0.1, lora_r: 8, SEQ_CLS as the task type and modules_to_save=["score"] in order to update and save the weights of the classification head we added to the model.
A.3 ICL
@@
We select 200 pairs for training and 60 pairs for development. To ensure balanced evaluation, we drew a random subset of examples, ensuring an equal number of labeled examples across the different relations. Within our DSPy program we incorporate chain-of-thought (CoT) reasoning and BootstrapFewShotWithRandomSearch optimizer.121212https://dspy-docs.vercel.app/docs/building-blocks/optimizers In CoT, a model is instructed to reason step by step to get to its answer, which can improve performance Wei et al. (2022). BootstrapFewShotWithRandomSearch is an enhanced version of the BootstrapFewShot teleprompter,131313https://dspy-docs.vercel.app/docs/deep-dive/teleprompter/bootstrap-fewshot which uses a teacher (GPT4 OpenAI (2023)) - student (GPT4o-mini) approach to generate few-shot demonstrations by sampling from the training data. During the training phase, the teacher model simulates potential outputs and rationales for labeled instances. The teacher model iteratively generates few-shot examples by attempting to predict the correct outputs based on the training data, learning from its own successes and refining also the generated rationale over time. This process is repeated multiple times, with a random search conducted over the generated demonstrations. The chosen program is the one that achieves the highest performance on the dev set. During each iteration on the test set, the program receives few-shot examples along with their bootstrapped rationales, assisting it in forming better rationales, thereby resulting in improved predictions.
A.4 Additional Ablation study
Relational definitions Ablation
To further examine the effect of relational definitions, we conducted an ablation study restricted to the Hard-10 subset (Table 9). Generating relational definitions for all mention pairs across the full SCICO test set is computationally infeasible, as it would require producing hundreds of thousands of pairwise definitions, therefore we focused on the smallest subset. Interestingly, while coreference scores improved only modestly, the hierarchy metrics saw larger gains, strengthening our hypothesis that relational definitions particularly help the model capture hierarchical structure. To balance feasibility and performance, we set the candidate selection threshold at 25%, which we found to offer a practical trade-off between computational cost and accuracy.
A.5 Statistical Significance Analysis
Statistical Significance
We evaluate whether improvements of our SciCo-RadarRelational model over the strongest baseline (SciCo-Mistral-7B) are statistically reliable using one-tailed paired bootstrap tests over topics.
Method
For each subset (Hard-10, Hard-20) and metric, we compute the observed difference and generate bootstrap samples by resampling topics with replacement. For each sample, we recompute the metric and record the resampled difference . The one-tailed -value is the proportion of samples in which the relational model does not outperform the baseline.
Results
On the Hard-10 subset, improvements across all hierarchy metrics and CoNLL F1 are statistically significant (). On the Hard-20 subset, the relational model yields statistically significant improvements in both Hierarchy F1 () and Hierarchy F1-50% (), while gains in Path Ratio and CoNLL F1 remain positive but are not statistically significant. These patterns are consistent with our hypothesis that relational definitions most strongly benefit the most challenging cases.
| Subset | Metric | ||
| Hard-10 | Path Ratio | +4.30 | 0.047 |
| Hard-10 | Hier F1-50% | +8.57 | 0.044 |
| Hard-10 | Hierarchy F1 | +6.72 | 0.047 |
| Hard-20 | Hierarchy F1 | +4.77 | 0.050 |
| Hard-20 | Hier F1-50% | +4.99 | 0.048 |
| Hard-20 | Path Ratio | +0.48 | 0.113 |
| Subset | Model | F1 | ||
| Hard-10 | Relational | 71.89 | +1.91 | 0.040 |
| Hard-20 | Relational | 76.23 | +0.64 | 0.086 |
| Model | Coreference | Hierarchy | Path | |
| CoNLL F1 | F1 | F1-50% | Ratio | |
| SciCo-RadarRelational def 25% | 71.90 | 56.98 | 45.93 | 53.85 |
| SciCo-RadarRelational def 60% | 71.96 | 57.71 | 46.25 | 54.31 |
| SciCo-RadarRelational def 100% | 72.33 | 58.15 | 46.82 | 55.17 |